# Grid Monitoring and Protection with Continuous Point-on-Wave Measurements and Generative AI

### v1

@@ 4.2 Performance Evaluations
We compared ISFD with two overcurrent protection baselines: (i) the “conventional” over-current relay protection with predetermined a fixed pickup current and inverse time function defined by the IEEE 37112 standard [63], and (ii) AOCR—an adaptive method proposed in [64]–with pickup current as a linear function of $10$-second moving-average current and minimum fault current in the zone.
All three methods required detection threshold settings. The ISFD threshold can be obtained analytically for a given false positive rate (FPR). In contrast, the threshold settings for the conventional and AOCR methods were obtained empirically through simulations to meet the FPR target.
@@
In the experiments, ISFD was implemented based on a $4$th-order Neyman’s smooth test with the doubling search procedure as described in Algorithm 1. We chose $C=42.5$ and $\lambda=20$. We generated $10,000$ samples under the anomaly-free condition for innovation autoencoder training. We adopted Adam optimizer with $lr=1e-5$ to train the innovation autoencoder. The threshold of the conventional method and AOCR are chosen manually to cap the FPR at $0.05$. The inverse-time characteristics of these two over-current relay-based methods are chosen according to the moderately inverse characteristics in IEEE standard [63].
@@
We evaluate the performance of each protection scheme by True Positive Rate (TPR), False Positive Rate (FPR), and detection delays. For each fault scenario, we conducted $1,000$ independent runs of the same system, driving the SDG current profile with $1,000$ different trajectories learned from the EPFL dataset. The trajectories were learned through a weaker version of innovation autoencoder (WIAE) with the $l_{2}$ distance between $(X_{t})$ and $(\hat{X}_{t})$ in the learning objective replaced by the Wasserstein distance. WIAE was trained by the EPFL current profile, and SDG current trajectories from the same distribution as the EPFL current profile can then be drawn as the output sequence $(\hat{X}_{t})$ of the decoder. The detailed implementation of WIAE can be found on Github [^2] https://github.com/Lambelle/WIAE. All metrics were calculated through $1,000$ Monte Carlo runs.
Results
@@
R4 (sympathetic tripping) ISFD 0.0779 0.0453 N/A
@@
R4 (sympathetic tripping) ISFD 0.0593 0.0257 N/A
@@
True Positive Rate of Detection: From the TPR column evaluated under faults F1, F2, and F3, respectively, the primary and backup relays should trip for all faults. We see that ISFD achieved the best detection accuracy for the primary and backup relays whenever a relay served as either a primary or a backup under all three faults. For example, under F1, the fault detection rate is $100\%$ at the primary relay R3 and $99.91\%$ at the backup relay R2. Similar performance were observed under F2 and F3. ISFD did have sympathetic tripping due to false positive detection at Relay R4 $7.8\%$ of the time due to the presence of SDG.
Both AOCR and the conventional did not perform well at the primary relay R3 under all three faults with detection rates under $90\%$. At the backup relays, AOCR and the conventional performed well under F1 but insufficient under F2 and F3. Sympathetic tripping occurred $100\%$ of the time under F1 and F3.
False Positive Rate of Detection: All methods achieved FPR smaller than $5\%$, since we computed their corresponding decision threshold by setting their empirical FPRs to be $0.05$, analytically or empirically. In particular, for ISFD, the decision threshold is computed analytically, and its FPR under all cases is smaller than $5\%$, showing that the innovation autoencoder transformer current measurements to IID uniform samples effectively.
Detection Delay: The delay column shows how quickly the detection decisions were rendered by the three methods. ISFD significantly outperformed both AOCR and the conventional, reducing the detection time by 90% at primary relays. At backup relays, ISFD performed better for most cases except under F1 (worse than AOCR) and under F2 (comparable with the conventional).
@@
Test statistics under ISFD: For ISFD, we plotted the histogram of the innovation sequence associated with the current CPOW measurements under three faults in Fig. 8. The top row of the figure shows the histograms of the latent processes under three faults. The bottom row shows the scatter plots of ISFD’s decision statistics.
The leftmost panel shows the case of F1 at R4 unrelated to F1. For this case, the test should be negative had there been no SDG. The top-left and top-right figures show that the histograms of the latent processes were approximately uniform, indicating that the observations were deemed normal. The scatter plots below show that the Neyman smooth test statistics were below the threshold most of the time. Thus the rates of positive tests was about $7\%$ and $5\%$, respectively.
The middle panel shows the case of F2 at R5. As the primary relay, R5 should trip with low detection delays. The top-middle figure shows that the distributions of innovations under fault and no-fault cases exhibit clear distinctions. The anomaly-free histogram was uniform, whereas the anomalous one showed significant deviation from the uniform distribution. Consequently, Neyman smooth test statistics for the anomalous case were mostly above the decision threshold, resulting in a TPR higher than $98\%$.
Test statistics under the conventional method: For the conventional method, the test statistic was based on the pickup current. We plotted the maximum current magnitude in each decision block with decision threshold $I_{pickup}$, shown in Fig. 9.
@@
Test statistics under AOCR: AOCR compares current maximum magnitude with an adaptively chosen threshold. We plotted the difference between the maximum current magnitude and the adaptively chosen thresholds in each decision block in Fig. 9. A difference greater than 00 implies fault for AOCR.
@@
The decision statistics shown in the leftmost panel of Fig. 10 was obtained under the case F1 at R4, where SDG induced an increase of fault current at R4, who shouldn’t trip during fault. Though AOCR computes the decision threshold adaptively, it is still comparing pre-fault and post-fault current magnitudes at R4. The statistics show that the discrepancy between anomaly-free and anomalous cases is smaller using AOCR rather than the conventional method, but the current magnitude increase’s negative impact on decision statistics persists. For $100\%$ of its time, the current magnitude at R4 exceeds the adaptive threshold computed based on a 10-second moving average of pre-fault current.
@@
The middle panel displays the current magnitude under the case of F2 at R5, where R5’s geographical closeness to the SDG introduces great randomness to the no-fault current. The fault current witnesses a decrease in magnitude compared to the no-SDG counterpart, leading to inaccurate detection results. Though AOCR tries to mitigate the problem by choosing the threshold adaptively, it cannot accommodate high volatility in SGD’s output power, thus resulting in lower detection accuracy.
@@ 5 Conclusion
The future power grid will be dominated by power electronics-operated energy resources with fast dynamics and rich stochasticity. The state-of-the-art grid monitoring and control is ill-equipped to address formidable challenges brought by large-scale integration of renewables, multi-sector electrification, and climate-change-induced natural disasters. Therefore, next-generation monitoring must be based on synchronized CPOW measurements to provide high-fidelity and high-resolution situational awareness and control. To this end, this paper offers an approach to compression and interactive networking of CPOW data streaming integrated with data analytics at local IED and control centers.
@@ Acknowledgement
The authors also wish to acknowledge the contribution of Mei-Jen Lee who performed some of the simulations.
This work was supported in part by the National Science Foundation under Grant EECS-2218110.

### v2

@@ 4.3 Baseline Comparisons
We compared the proposed foundation model adaptation for over-current protection with two protection baselines: (i) the “conventional” over-current relay protection with a predetermined pickup current and the inverse time function defined by the IEEE 37112 standard [70], and (ii) AOCR—an adaptive method proposed in [71]–with pickup current as a linear function of $10$-second moving-average current and minimum fault current in the protection zone. All three methods required setting detection thresholds. The SNST threshold was obtained analytically for a given false positive rate (FPR). The threshold settings for the conventional and AOCR methods were obtained empirically through simulations to meet the FPR target.
@@
For the proposed foundation model in the experiments, the learning of the IRM autoencoder pair and the discriminator in Fig 3 are convolutional neural networks with three layers. The filter size is set to 20, and the first layer’s input dimension is 50. Each layer’s output dimensions are 100, 50, and 25. Batch normalization is applied to the input layer. The Adam optimizer is used to get the optimal weights for all neural network structures. The SNST used in the foundation model adaptation was implemented based on a $4$th-order Neyman’s smooth test with the doubling search procedure as described in Algorithm 1. We chose $C=42.5$ and $\lambda=20$. We generated $10,000$ samples under the anomaly-free condition for innovation autoencoder training. We adopted the Adam optimizer with $lr=1e-5$ to train the innovation autoencoder. The threshold of the conventional method and AOCR are chosen manually to cap the FPR at $0.05$. The inverse-time characteristics of these two over-current relay-based methods are chosen according to the moderately inverse characteristics in IEEE standard [70].
@@
We evaluated the performance of each protection scheme by the True Positive Rate (TPR), the False Positive Rate (FPR), and the detection delay. For each fault scenario, we conducted $1,000$ independent runs of the same system, driving the SDG current profile with $1,000$ different trajectories learned from the EPFL dataset.
The trajectories were learned through a weaker version of innovation autoencoder (WIAE) with the $l_{2}$ distance between $(X_{t})$ and $(\hat{X}_{t})$ in the learning objective replaced by the Wasserstein distance. WIAE was trained by the EPFL current profile, and SDG current trajectories from the same distribution as the EPFL current profile can then be drawn as the output sequence $(\hat{X}_{t})$ of the decoder. The detailed implementation of WIAE can be found on Github [^2] https://github.com/Lambelle/WIAE. All metrics were calculated through $1,000$ Monte Carlo runs.
We evaluated the performance of three candidate solutions using the IEEE 13-bus distribution network shown in Fig. 6 (left), where multiple fault scenarios were considered in our study. To account for randomness in the measurements introduced by renewable energy resources, we added a $1$ MW SDG controlled by a real-world current profile collected at the micro-grid located in EPFL between node $692$ and $675$. The SDG was driven by EPFL’s current profile at the battery connection [^3] Data can be found here: https://github.com/DESL-EPFL/Point-on-wave-Data-of-EPFL-campus-Distribution-Network, and current measurement samples were collected at $50$ KHz sampling frequency. The entire system is simulated through MATLAB/Simulink.
4.4 Numerical Results
When a single-phase short circuit fault happens at F1, R3 is the primary relay for F1 and R2 as a backup. Sympathetic tripping at R4 may occur under F1 because the current feed from SDG downstream could increase significantly. Similarly, R4 is also subject to sympathetic tripping under F3 due to current flow from downstream DG. In this case, the primary relay is R2, and the backup is R1. On the other hand, the primary relay R5 for fault scenario F2 faces protection blinding because of DG’s support for current and voltage at R5 during fault at F2.
@@ 4.4 Numerical Results
R4 (sympathetic tripping) SNST 0.0779 0.0453 N/A
@@
R4 (sympathetic tripping) SNST 0.0593 0.0257 N/A
@@
True Positive Rate Performance: From the TPR column evaluated under faults F1, F2, and F3, respectively, the primary and backup relays should trip for all faults. We see that SNST achieved the best detection accuracy for the primary and backup relays whenever a relay served as either a primary or a backup under all three faults. For example, under F1, the fault detection rate is $100\%$ at the primary relay R3 and $99.91\%$ at the backup relay R2. Similar performance were observed under F2 and F3. SNST did have sympathetic tripping due to false positive detection at Relay R4 $7.8\%$ of the time due to the presence of SDG. SNST improved TPR in primary relay by 5.7%, 106.9 %, and 67.3% under F1, F2 and F3, respectively.
Both AOCR and the conventional did not perform well at the primary relay under F2 and F3, with detection rates under $90\%$. At the backup relays, AOCR and the conventional performed well under F1 but insufficient under F2 and F3. Sympathetic tripping occurred $100\%$ of the time under F1 and F3.
False Positive Rate of Detection: All methods achieved FPR smaller than $5\%$, since we computed their corresponding decision threshold by setting their empirical FPRs to be $0.05$, analytically or empirically. In particular, for SNST, the decision threshold is computed analytically, and its FPR under all cases is smaller than $5\%$, showing that the innovation autoencoder transformer current measurements to IID uniform samples effectively.
Detection Delay Performance: The delay column shows how quickly the detection decisions were rendered by the three methods. Detection delay is defined as the latency between fault time and fault detection time. SNST significantly outperformed both AOCR and the conventional, reducing the detection time by 90% at primary relays. At backup relays, SNST performed better for most cases except under F1 (worse than AOCR) and under F2 (comparable with the conventional).
@@
Insights from test statistics for the SNST detector: For the proposed SNST detection, we plotted the histogram of the innovation sequence associated with the current WMU measurements under three faults in Fig. 7. The top row of the figure shows the histograms of the latent processes under three faults. The bottom row shows the scatter plots of SNST’s decision statistics.
The leftmost panel shows the case of F1 at R4 unrelated to F1. For this case, the test should have been negative had there been no SDG. The top-left and top-right figures show that the histograms of the latent processes were approximately uniform, indicating that the observations were deemed normal. The scatter plots below show that the Neyman smooth test statistics were below the threshold most of the time. Thus, the rates of positive tests were about $7\%$ and $5\%$, respectively.
The middle panel shows the case of F2 at R5. As the primary relay, R5 should trip with low detection delays. The top-middle figure shows that the distributions of innovations under fault and no-fault cases exhibit clear distinctions. The anomaly-free histogram was uniform, whereas the anomalous one showed a significant deviation from the uniform distribution. Consequently, Neyman smooth test statistics for the anomalous case were mostly above the decision threshold, resulting in a TPR higher than $98\%$.
Test statistics for the conventional detector: For the conventional method, the test statistic was based on the pickup current. We plotted the maximum current magnitude in each decision block with decision threshold $I_{pickup}$, shown in Fig. 8.
@@
Test statistics for the AOCR detector: AOCR compared the maximum current magnitude with an adaptively chosen threshold. We plotted the difference between the maximum current magnitude and the adaptively chosen thresholds in each decision block in Fig. 9. A difference greater than 00 implies fault for AOCR.
@@
The decision statistics shown in the leftmost panel of Fig. 9 were obtained under the case F1 at R4, where SDG induced an increase of fault current at R4, which should not have tripped during fault. Though AOCR computed the decision threshold adaptively, it still compared the pre-fault and post-fault current magnitudes at R4. The statistics showed that the discrepancy between anomaly-free and anomalous cases was smaller using AOCR rather than the conventional method, but the negative impact of the current magnitude increase on decision statistics persisted. For $100\%$ of its time, the current magnitude at R4 exceeded the adaptive threshold computed based on a 10-second moving average of pre-fault current.
@@
The middle panel displays the current magnitude under the case of F2 at R5, where R5’s geographical closeness to the SDG introduced significant randomness to the no-fault current. The fault current witnessed a decrease in magnitude compared to the no-SDG counterpart, leading to inaccurate detection results. Though AOCR tried to mitigate the problem by choosing the threshold adaptively, it could not accommodate the high volatility in SGD’s output power, resulting in lower detection accuracy.
5 Adaptation for Synchro-waveform Compression
This section presents the foundation model adaptation to compression of synchro-waveform streaming with the goal of facilitating communications by allocating the same or slightly higher communication bandwidth required by PMU data streaming systems. The raw sampling rates of WMU data are one to three orders of magnitude higher than PMU, as shown in Fig. 1. To capture some of the IBR-related events at the same bandwidth as required by PMU systems, WMU data streaming needs to have compression ratios of 10-100:1 while maintaining the same level of decompression accuracy.
We propose here a foundation model adaptation for lossy WMU data compression, built on the subband coding idea first proposed in [6]. Instead of compressing raw WMU streaming, we propose compressing and communicating the innovation sequence extracted by the IRM, which is the data feature sufficient for monitoring, control, and real-time decision-making.
Figure 10: A schematic of the foundation model adaptation for WMU data compression.
The schematic of the proposed compression system is shown in Fig. 10 with key components explained next. The upper branch of Fig. 10 is the foundation model adaptation schematic in the intelligent electronic devices (IEDs) where current and voltage WMU data are generated. The lower branch resides in the substations or control centers where monitoring and control decisions are made. Power system signals are narrow passband signals with special subband spectral structure illustrated in Fig. 10. Around the fundamental frequency ±$\pm f_{0}$ is the fundamental frequency band regulated with frequency deviations within $W_{0}$. Typically $W_{0}<1$ Hz. Above the fundamental frequency band are harmonic subbands at the multiples of $f_{0}$, $f_{k}=kf_{0}$, that account for higher-order harmonics due to nonlinear devices in the network. There may also be inter-harmonic frequency components represented by the grey area. Higher-order harmonics and inter-harmonics may appear sporadically; they are modeled as on-off random processes.
Subband Coding
The A/D converter samples the WMU $x(t)$ at the Nyquist rate to obtain a digital continuous point-on-wave sequence $x[n]$. Depending on the spectral bandwidth of $x(t)$, the Nyquist sampling rate is typically in the tens of kHz (or higher) range. The subband coding scheme proposed in [6] greatly reduces the Nyquist-sampled data rate via the subband decomposition that encodes each harmonic (and inter-harmonic) subband separately. By the passband Nyquist sampling theorem, a harmonic subband with passband bandwidth $W$ requires only a $2W$ sampling rate. For instance, if the fundamental subband has the bandwidth of $W_{0}=1$ (Hz), two samples per second or one sample per 30 cycles is sufficient to capture the signal in the fundamental frequency band fully. For an $m$ harmonic-subband spectrum, each with $W$ Hz passband bandwidth, $mW/30$ samples per cycle are sufficient for lossless compression. The subband coding block extracts the baseband components of $X(f)$, down-sampling the Nyquist-sampled data to achieve two to three orders of magnitude data rate reduction. The subband coding technique is well developed for multi-media compression and is embedded in practical audio/video communications data-streaming technologies. The subband encoder proposed in [6] can be easily implemented by off-the-shelf filter banks. See details in [6]. Note the per-cycle data samples of the subband WMU signals $\{(x_{i}[n])\}$ are baseband representations of the signal spectrum, typically at the rate of less than one sample per cycle under normal operating conditions.
Innovation Autoencoder and Local Data Analytics
While the subband coding scheme significantly reduces the data rate of the direct Nyquist sampled $(x[n])$, further data rate reductions can be achieved through information-theoretic approaches that offer an optimal tradeoff between the compression rate and distortions of reconstruction. Additional data rate reduction can be achieved through local data analytics that adaptively select harmonic subbands for transmission. For instance, harmonic and inter-harmonic subbands may be inactive under normal operating conditions. During significant network events, signal contents from only a few harmonic subbands may be informative.
The innovation autoencoder of the pre-trained foundation model discussed in Sec. 2 is used to transform subband signals $\{(x_{i}[n])\}$ to a set of innovations sequences $\{(v_{i}[n])\}$, each is an IID uniform or Gaussian sequences. For optimal data compression over the typical additive Gaussian noise channel of the utility network with innovation sequence $(v_{i}[n])$, the optimal bits-allocation to each subband innovation can be determined. This is discussed as part of the vector quantization encoder in Sec. 5.
The local data analytics unit detects anomalies and novelties using techniques presented in Sec. 3. The local data analytics unit produces the local state sequence $(s_{t})$ that can be used to compress data further and provide summary statistics for the control center. Specifically, $(s_{t})$ may include the normal/fault state sequence and estimated fault locations when faults are detected locally. Under the normal system state, data from many subbands can be suppressed, and the control center can recall the suppressed signal if necessary.
Vector Quantization Encoder
Vector quantization converts innovations and state sequence $\{(\nu_{i}[n]),s[n])\}$ to bit-streams to be transmitted over the utility network (after multiplexing with bit streams from other IEDs.) The quantization schemes developed in [6] can be used. The new feature considered here is an information theoretic approach to compression not included in [6].
Consider the IID-Gaussian innovation sequences $(\nu_{i}[n])$, with mean and variance $(\mu_{i},\sigma_{i}^{2})$ for the $i$th subband innovations. Given the acceptable mean-square distortion of the reconstruction $D_{i}:=\mathbb{E}(\hat{x}_{i}[n]-x_{i}[n])^{2}$, the optimal compression is achieved by allocating distortion $D_{i}$ to subband $i$ constrained by $D=\sum_{i}D_{i}$ through the so-called inverse water-filling procedure [72]. The Gaussian rate-distortion function gives the minimum transmission rate subject to the distortion measure
$R(D)=\sum_{i=1}^{m}\frac{1}{2}\log\frac{\sigma_{i}^{2}}{D_{i}}.$
Decoder at Substations or Control Center
The lower branch of Fig. 10 describes the building blocks of the decoding process at the control center. For the most part, this process is the inverse of the upper branch, following mostly the procedure described in [6]. The vector quantization decoder reproduces the innovation sequence $\{(\hat{\nu}_{i}[n])\}$ for each subband (for those subbands that are not suppressed.) The innovation autodecoder is used to produce the decompressed subband signal $\{(\hat{x}_{i}[n])\}$ (in the baseband form), and the subband decoder assembles the subband signal to form $\hat{x}[n]$ as the estimation of the original WMU data $x[n]$, for which the estimation error is controlled by $D$ in the rate-distortion function.
Central Data Analytics and Interactive Networking
Integrating data analytics with compression and networking provides the key component for interactive networking: the local IEDs can initiate the interaction based on local detection and transmit anomalous waveforms. Likewise, with the received data feeds from multiple IEDs, the central data analytics unit can initiate requests for high-resolution data when data feeds result in ambiguities and inference with low confidence. Such types of “push-pull” networking protocols have been widely used in streaming, Internet-of-Things, and content delivery networking technology [11].
@@ 6 Conclusion
The future power grid will be dominated by power electronics-operated energy resources with fast dynamics and rich stochasticity. The existing grid monitoring and control framework is inadequate to address new challenges brought by large-scale integration of renewables, multi-sector electrification, and climate-change-induced natural disasters, and the next-generation monitoring and control need to include high-resolution synchro-waveform measurement technology and leverage powerful AI tools.
The proposed physics-based AI foundation model is an initial attempt to extend the successful LLM-based foundation models to meet challenges in power system operations. Among the many challenges of this endeavor that have yet to be addressed here is the complexity involved in pre-training foundation models for power system applications. Unlike the LLM-based foundation model, where a single foundation model can be used for all text-based applications of the same language, a foundation model for power systems needs to be locational, as measurement at each location has its own characteristics and constraints. Although the innovation-based foundation model is considerably smaller than the LLM-based foundation model, the amount of pre-training can be significant. Using transfer learning techniques, as we have observed in our experiments, can significantly reduce the complexity of pre-training.
@@ Acknowledgement
The authors also wish to acknowledge the contribution of Mei-Jen Lee who performed some of the simulations. We also thank the anonymous reviewers whose comments helped to improve the presentation of this work.
This work was supported in part by the National Science Foundation under Awards 2218110 and 2419622.
The authors have no relevant financial or non-financial interests to disclose.